Consider the following recent financials for XYZ Corporation:

Income Statement

Balance Sheet

Sales

73,802

Assets

209,087

Debt

38,278

Costs

44,281

Equity

170,809

EBIT

29,521

Taxes @ 38%

11,218

Total

209,087

Total

209,087

Net Income

18,303

    

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2,907 was paid, and the company wishes to maintain a constant payout ratio. Next year’s sales are projected to grow by 25%.

What is the pro-forma value for equity? (Round answer to 2 decimal places. Do not round intermediate calculations).

What is the external financing needed using the pro-forma approach? (Round answer to 2 decimal places. Do not round intermediate calculations).
Answer #1

Current Sales = $ 73802, Expected Sales Growth = 25 % and Expected Sales = 73802 x 1.25 = $ 92252.5

Expected Costs = 44281 x 1.25 = $ 55351.25 (as cost is proportional to sales)

Expected EBIT = 92252.5 - 55351.25 = $ 36901.25

Expected Taxes @ 38 % = 0.38 x 36901.25 = $ 14022.5

Expected Net Income = 36901.25 - 14022.5 = $ 22878.75

Current Net Income = $ 18303 and Dividends = $ 2907

Current Payout Ratio = 2907 / 18303 ~ 0.1588

Expected Dividend (if the same dividend payout ratio is maintained) = 0.1588 x 22878.75 ~ $3633.75

Retained Earnings (Expected) = Expected Net Income - Expected Dividend = 22878.75 - 3633.75 ~ $ 19245

Pro-Forma Value of Equity = Current Value of Equity + Expected Retained Earnings = 170809 + 19245 = $ 190054

Expected Value of Assets = 209087 x 1.25 ~ $ 261359

External Financing Needed = Expected Value of Assets - Debt - Equity - Retained Earnings = 261359 - 38278 - 170809 - 19245 ~ $ 33026.75